an automobile engine can run on a mixture of gasoline and a substitute fuel. if gas costs $3.25 per gallon and the substitute fuel costs $2 per gallon, what percent of mixture must be substitute fuel to bring the cost down to $3.00 per gallon?
Answer by stanbon(75887)

an automobile engine can run on a mixture of gasoline and a substitute fuel. if gas costs $3.25 per gallon and the substitute fuel costs $2 per gallon, what percent of mixture must be substitute fuel to bring the cost down to $3.00 per gallon?
--------------------
Equation for on2 gallon of mixture::
3.25x + 2(1-x) = 3.00
----
325x + 200 - 200x = 300
125x = 100
x = 0.8 gallons (amt of $3.25 gasoline)
1-x = 0.2 gallons (amt of $2.00 gasoline)
-----
Ans: 0.2 = 20% substitute fuel
----------------
Cheers,
Stan H.
------------------